A process has a mean of 9 grams and a standard deviation of 0.25 grams.  The lower spec limit is 8.00 grams and the upper spec limit is 10.00 grams.

- Compute Cp.

- What is the probability the process will operate within the specification limits?

- P{out of spec} for Wonderdisk

- μ = 10, σ = 2.3

qP{X ≤ 6 or X ≥ 12}

qP{8 ≤ X ≤ 13}
Answer #1

Answer:

PAGE NO. DATE 6 (an) The process has a mean of a grams and a standard deviation of 0.25 gramu. 7 Given:- 439_qramy z XNN (9,0

PAGE NO // The values can be found wring Stanishal aber Plout of specification) = |- P ( within specification limit =1 -0.999

PAGEHO PC:85x5133 =P(8-10 SX510 13-10 2.3. 2.3 2.3 = P(-0.869565€ Z 41.30434) = P(1.30434) - P(-0.869565) = 0.90394 0.19227 1

NOTE:: I HOPE YOUR HAPPY WITH MY ANSWER....***PLEASE SUPPORT ME WITH YOUR RATING...

***PLEASE GIVE ME "LIKE"...ITS VERY IMPORTANT FOR ME NOW....PLEASE SUPPORT ME ....THANK YOU